# A partially broken Fronsdal model for massless higher-spin particles of integer spin

### v2

@@ 4.2 Equations of motion and degrees of freedom
The equations of motion corresponding to the PBF model ${\pazocal L}^{(s)}(1,1,c)$, see (40), correspond to the vanishing of the rank-s double-traceless Euler tensor:
$E=\Box h-\partial\ \partial\!\cdot\!h+\eta\ \partial\!\cdot\!\partial\!\cdot\!h+\partial\partial h^{\prime}-\eta\Box h^{\prime}-\frac{c}{2}\eta\ \partial\ \partial\!\cdot\!h^{\prime}-\Upsilon\ \eta\ \eta\ \partial\!\cdot\!\partial\!\cdot\!h^{\prime}=0,$ (48)
@@
$\Upsilon=\frac{(1-c)[D+2s-6]}{(D+2)[D+4(s-4)]}.$ (49)
The equations (48) are invariant under trDiff:
$\delta h=\partial\,\overline{\Lambda}^{\ddagger}\quad.$ (50)
In agreement with the gauge symmetry (50), the divergence of the equations of motion is a linear combination of $\eta$ and two derivatives $\partial\,\partial$,
$\partial\!\cdot\!E=\eta\left[\partial^{3}\!\cdot\!h-\left(1+\frac{c}{2}\right)\Box\partial\!\cdot\!h^{\prime}-\left(\Upsilon+\frac{c}{2}\right)\partial\ \partial\!\cdot\!\partial\!\cdot\!h^{\prime}-\Upsilon\eta\ \partial^{3}\!\cdot\!h^{\prime}\right]+(1-c)\partial\ \partial\ \partial\!\cdot\!h^{\prime}=0.\\$ (51)
@@
$\partial\!\cdot\!h^{\prime}$ $=$ $0,$ (52)
$\partial\!\cdot\!\partial\!\cdot\!\partial\!\cdot\!h$ $=$ $0.$ (53)
With those constraints, the trace of the equations of motion imply $(\Box h^{\prime}-\partial\!\cdot\!\partial\!\cdot\!h)=0$. Back in (48) we obtain the vanishing of the Fronsdal tensor as in the Fronsdal model, i.e.,
${\pazocal F}=\Box h-\partial\ \partial\!\cdot\!h+\partial\ \partial\ h^{\prime}=0.$ (54)
All the equations (48)-(54) are invariant under the gauge transformations (50). Now we fix the spin-$s$ de Donder gauge,
$\overline{f}=\partial\cdot h-\frac{1}{2}\partial\,h^{{}^{\prime}}=0.$ (55)
The number of gauge conditions may not exceed the number of independent gauge parameters. Because of the gauge invariant constraints (52) and (53), the traceless rank-(s-1) symmetric tensor (55) has vanishing double divergence on shell: $\partial\!\cdot\!\partial\!\cdot\!\overline{f}=0$. Thus, we are effectively fixing on shell the same number of gauge conditions as the number of independent gauge parameters $\overline{\Lambda}^{\ddagger}$. Using the de Donder gauge in (54) we obtain the massless equations $\Box h=0$.
All the equations so far, including the gauge conditions (55) are invariant under residual harmonic ( $\Box\overline{\Lambda}^{\ddagger}=0$) gauge transformations which allows us to “regauge”. Now we proceed with the counting of degrees of freedom (d.o.f.). The double-traceless field $h$ has
$N_{h}=\binom{D+s-1}{D-1}-\binom{D+s-5}{D-1}$ (56)
@@
$N_{\Lambda}=\binom{D+s-2}{D-1}-\binom{D+s-4}{D-1}-\left[\binom{D+s-4}{D-1}-\binom{D+s-6}{D-1}\right$]. (57)
The equations of motion lead to the gauge independent constraints (52) and (53) which are traceless rank-(s-3) symmetric tensors. The number of independent constraints is
$N_{\text{constraints}}=2\left[\binom{D+s-4}{D-1}-\binom{D+s-6}{D-1}\right$] (58)
@@
$N_{dof}$ $=$ $N_{h}-2N_{\Lambda}-N_{\text{constraints}}$ (59)
@@
For arbitrary $D$ dimensions and arbitrary spin-$s$, (59) coincides exactly with the number of d.o.f. of the Fronsdal model. The inclusion of higher derivative Lagrangian constraints in the counting of degrees of freedom is typical of gauge parameters with derivative restrictions, see chapter 5 of [8]. This is the case also of the spin-2 TDiff model (24) whose equations of motion lead to the two derivative scalar constraint $\partial^{\alpha}\partial^{\beta}h_{\alpha\beta}=0$. So, although we have (in $D=4$) only 3 independent gauge parameters due to the transverse condition $\partial^{\mu}\epsilon^{T}_{\mu}=0$, we can fix the four gauge conditions $f_{\mu}=\partial^{\mu}h_{\mu\nu}=0$ since they are on shell transverse due to the two derivative scalar constraint. Taking into account the “regauging” we have $N_{dof}=10-2\times 3-1=3$ in agreement with one spin-0 and the $\pm 2$ helicities of the spin-2 particle [^8] Remarkably, the Hamiltoninan analysis of [19] has shown that the spin-2 TDiff model has 2D12D-1, instead of the expected 2(D1)2(D-1), first class constraints and no second class constraints. The Hamiltonian constraints analysis of the PBF model for arbitrary integer spin-$s$ is in progress..
@@ 4.3.1 The spin-3 case
$\delta h_{\mu\nu\alpha}=ip_{(\mu}\overline{\Lambda}^{\ddagger}_{\nu\alpha)}.$ (60)
@@
$p\!\cdot\!p\!\cdot\!\overline{\Lambda}^{\ddagger}=0=\eta\cdot\overline{\Lambda}^{\ddagger}.$ (61)
Assuming $p_{+}\neq 0$, due to (61), the gauge parameters $\overline{\Lambda}^{\ddagger}_{--}$ and $\overline{\Lambda}^{\ddagger}_{+-}$ become functions of $(\overline{\Lambda}^{\ddagger}_{++},\overline{\Lambda}^{\ddagger}_{IJ})$. However, we can use the whole set of unconstrained parameters $(\overline{\Lambda}^{\ddagger}_{++},\overline{\Lambda}^{\ddagger}_{IJ},\overline{\Lambda}^{\ddagger}_{+J},\overline{\Lambda}^{\ddagger}_{-J})$ in order to impose the maximal number of gauge conditions:
$h_{+++}=0=h_{+IJ}=h_{++J}=h_{+-J}\,.$ (62)
@@
$E_{\mu\nu\rho}(p)\equiv p^{2}h_{\mu\nu\alpha}-p_{(\mu}p^{\rho}h_{\nu\alpha)\rho}+p_{(\mu}p_{\nu}h_{\alpha)}=0.$ (63)
Henceforth we use the short hand notation: $(\mu,\nu,\rho)\equiv E_{\mu\nu\rho}(p)$ and assume that (62) is holding true. From $(+,+,+)=0$ one finds $h_{-++}=0$. The equations $(+,+,J)=0$ and $(+,I,J)=0$ lead to $h_{IIJ}=0$ and $h_{-IJ}=\frac{1}{p^{+}}p_{K}h_{IJK}$, respectively, they lead altogether to $h^{-JJ}=0$. The constraint $\partial\!\cdot\!h^{\prime}=0$ implies $h_{+--}=0$. The field is thus on shell traceless ($h_{\alpha}=-2\,h_{+-\alpha}+h_{II\alpha}=0$) and satisfies $h_{+\mu\nu}=0$. From $(+,\nu,\alpha)=0$ we have the transverse constraint $p^{\rho}h_{\rho\nu\alpha}=0$ which at $(\nu,\alpha)=(-,k)$ and $(\nu,\alpha)=(-,-)$ leads, using previous results, to
$h_{--K}=\frac{p_{I}\,p_{J}\,h_{IJK}}{p_{+}^{2}}\quad;\quad h_{---}=\frac{p_{I}\,p_{J}\,p_{K}h_{IJK}}{p_{+}^{3}}.$ (64)
So we can consider $h_{IJK}$ as the only relevant field components. With $h_{\alpha}=0=p^{\rho}h_{\rho\mu\nu}$ back in (63) we have $p^{2}h_{\mu\nu\alpha}=0$. Effectively, we have the following physical equations:
$p^{2}h_{IJK}=0,\quad h_{JJK}=0.$ (65)
@@
$p^{2}h_{\mu\nu\alpha\beta}-p_{(\mu}\ p^{\rho}h_{\rho\nu\alpha\beta)}+p_{(\mu}\ p_{\nu}\ h^{\prime}_{\alpha\beta)}\equiv(\mu,\nu,\alpha,\beta)=0.$ (67)
@@
$h_{KKIJ}=0.$ (68)
From (+++)=0(-+++)=0 and the constraint $(p\cdot h^{{}^{\prime}})_{+}=0$, we have $h_{++--}=0=h_{+-KK}$. Now we make use of the only unconstrained gauge parameters left over. Namely, since $\delta h_{+-IJ}=i\,p_{(+}\overline{\Lambda}^{\ddagger}_{IJ-)}$ and we already have $h_{+-KK}=0$, we only need the space traceless parameter $\overline{\Lambda}^{\ddagger}_{\overline{I}\overline{J}-}=\overline{\Lambda}^{\ddagger}_{IJ-}-\delta_{IJ}\overline{\Lambda}^{\ddagger}_{KK-}/(D-2)$ in order to fix our last light cone gauge condition: $h_{+-\overline{I}\overline{J}}=0$. Consequently, we have $h_{+-IJ}=0$ henceforth. There is no remaining unconstrained parameter anymore, since $\overline{\Lambda}^{\ddagger}_{KK-}$ is already constrained as we have argued before (66).
@@
$h_{---I}=\frac{p_{J}\,p_{K}\,p_{L}\,h_{IJKL}}{p_{+}^{3}}\quad$ ; $h_{----}=\frac{p_{I}\,p_{J}\,p_{K}\,p_{L}\,h_{IJKL}}{p_{+}^{3}}$ (69)
@@
$p^{2}h_{IJKL}=0.$ (70)
Altogether with (68) we have (D+4)(D3)(D2)(D1)/(4!)(D+4)(D-3)(D-2)(D-1)/(4\ !) massless propagating modes as in the Fronsdal spin-4 model.
@@ 4.3.3 Arbitrary integer spin
$h_{+(k)J(s-k)}=0\quad,\quad k=1,\cdots,s\,.$ (71)
@@
$p^{2}\,h_{\mu_{1}\cdots\mu_{s}}+p_{(\mu_{1}}p_{\mu_{2}}\,h^{{}^{\prime}}_{\mu_{3}\,\mu_{4}\cdots\mu_{s})}-p_{(\mu_{1}}p^{\rho}h_{\rho\,\mu_{2}\cdots\mu_{s})}\equiv\left(\mu_{1}\,\mu_{2}\cdots\mu_{s}\right)=0.$ (72)
@@
$h_{++\mu(s-2)}=0\quad\,\text{and}\,\quad h_{+KK\mu(s-3)}=0,$ (74)
@@
$h^{-(l)I(s-l)}=\frac{1}{(p^{+})^{l}}p^{I_{1}}p^{I_{2}}\dots p^{I_{l}}h^{I_{1}I_{2}\dots I_{l}I(s-l)},$ (76)
where $l\in\{1,2,\dots,s\}$. Notice that the constraint $p\cdot p\cdot p\cdot h=0$ is equivalent to (76). So the only nontrivial field components are given by $h_{J(s)}$. The equations of motion $\left(I_{1}\cdots I_{s}\right)=0$ become simply
$p^{2}h^{I(s)}=0.$ (77)
Since we still have the algebraic constraints (73) we end up with two helicities in $D=4$. The condition on the gauge parameter $\partial\cdot\partial\cdot\overline{\Lambda}^{\ddagger}=0$ is compensated by the constraint $\partial\cdot h^{{}^{\prime}}=0$. Notice that they are both totally symmetric rank-(s-3) traceless tensors.
@@ 5 Conclusion and perspectives
We have found here a partially broken Fronsdal (PBF) model for arbitrary integer spin, it is defined in (47), (46) and (40). The model is invariant under traceless higher spin diffeomorphisms with vanishing double divergence, see (50). In subsection 3.3 we have verified the content of the spin-3 PBF model in a gauge invariant way while in subsection 4.3 we have checked the spectrum of the arbitrary integer spin case with $s\geq 3$ in the light cone gauge. We conclude that the PBF model has the same propagating content of the Fronsdal model, i.e., only the $\pm s$ helicities in $D=4$.
@@
$\partial\cdot T=\eta\,\Omega+\partial\,\partial\,J,$ (78)
where $\Omega$ and $J$ are rank-(s-3) symmetric tensors constrained only by the fact that the double trace of the right hand side of (78) must vanish for $s\geq 5$. In the Fronsdal theory the non conservation law is more restrictive ($J=0$).
@@
Regarding a curved background, since $\nabla\cdot h^{{}^{\prime}}$ is by itself invariant under $\delta h=\nabla\overline{\Lambda}^{\ddagger}$, with $g^{\mu_{1}\mu_{2}}\,\overline{\Lambda}^{\ddagger}_{\mu_{1}\mu_{2}\cdots\mu_{s-1}}=0=\nabla\cdot\nabla\cdot\overline{\Lambda}^{\ddagger}$, it follows that the same non minimal compensating terms of the Fronsdal action [25] on (A)dS backgrounds, quadratic on $h$ and $h^{{}^{\prime}}$, will turn the model $\pazocal{L}^{(s)}(1,1,c)$ given in (41), with $\partial_{\mu}\to\nabla_{\mu}$, invariant under gauge transformations on (A)dS.
@@
It is important to mention that the PBF model is a two parameters theory, $\pazocal{L}_{PBF}^{(s)}=\pazocal{L}^{(s)}[a,b_{S}^{(s)}(a,c),c]=\pazocal{L}_{PBF}^{(s)}(a,c)$. We can always redefine one of the parameters as we wish by means of local field redefinitions that we have named r-shifts, see (41) and (42). We are currently investigating the role of the remaining parameter which can not be shifted, along the lines of [21] where we have shown, inspired also by non-local field redefinitions, that the two parameters TDiff spin-3 model $S[a,b,c_{S}(a,b)]$, see (27), is physically equivalent to its simplest version $S(0,0,0)$ which corresponds to the spin-3 case of the Maxwell-like model of [22]. Similarly, we believe that non-local field redefinitions may help to clarify the role of the left over constant parameter of the PBF model.

### v3

@@ 4.2 Equations of motion and degrees of freedom
The equations of motion corresponding to the PBF model ${\pazocal L}^{(s)}(1,1,c)$, see (4.1), correspond to the vanishing of the rank-s double-traceless Euler tensor:
$E=\Box h-\partial\ \partial\!\cdot\!h+\eta\ \partial\!\cdot\!\partial\!\cdot\!h+\partial\partial h^{\prime}-\eta\Box h^{\prime}-\frac{c}{2}\eta\ \partial\ \partial\!\cdot\!h^{\prime}-\Upsilon\ \eta\ \eta\ \partial\!\cdot\!\partial\!\cdot\!h^{\prime}=0,$ (4.9)
@@
$\Upsilon=\frac{1-c}{D+2(s-4)}.$ (4.10)
The equations (4.9) are invariant under trDiff:
$\delta h=\partial\,\overline{\Lambda}^{\ddagger}\quad.$ (4.11)
In agreement with the gauge symmetry (4.11), the divergence of the equations of motion is a linear combination of $\eta$ and two derivatives $\partial\,\partial$,
$\partial\!\cdot\!E=\eta\left[\partial^{3}\!\cdot\!h-\left(1+\frac{c}{2}\right)\Box\partial\!\cdot\!h^{\prime}-\left(\Upsilon+\frac{c}{2}\right)\partial\ \partial\!\cdot\!\partial\!\cdot\!h^{\prime}-\Upsilon\eta\ \partial^{3}\!\cdot\!h^{\prime}\right]+(1-c)\partial\ \partial\ \partial\!\cdot\!h^{\prime}=0.\\$ (4.12)
@@
$\partial\!\cdot\!h^{\prime}$ $=$ $0,$ (4.13)
$\partial\!\cdot\!\partial\!\cdot\!\partial\!\cdot\!h$ $=$ $0.$ (4.14)
With those constraints, the trace of the equations of motion imply $(\Box h^{\prime}-\partial\!\cdot\!\partial\!\cdot\!h)=0$. Back in (4.9) we obtain the vanishing of the Fronsdal tensor as in the Fronsdal model, i.e.,
${\pazocal F}=\Box h-\partial\ \partial\!\cdot\!h+\partial\ \partial\ h^{\prime}=0.$ (4.15)
All the equations (4.9)-(4.15) are invariant under the gauge transformations (4.11). Now we fix the spin-$s$ de Donder gauge,
$\overline{f}=\partial\cdot h-\frac{1}{2}\partial\,h^{{}^{\prime}}=0.$ (4.16)
The number of gauge conditions may not exceed the number of independent gauge parameters. Because of the gauge invariant constraints (4.13) and (4.14), the traceless rank-(s-1) symmetric tensor (4.16) has vanishing double divergence on shell: $\partial\!\cdot\!\partial\!\cdot\!\overline{f}=0$. Thus, we are effectively fixing on shell the same number of gauge conditions as the number of independent gauge parameters $\overline{\Lambda}^{\ddagger}$. Using the de Donder gauge in (4.15) we obtain the massless equations $\Box h=0$.
All the equations so far, including the gauge conditions (4.16) are invariant under residual harmonic ( $\Box\overline{\Lambda}^{\ddagger}=0$) gauge transformations which allows us to “regauge”. Now we proceed with the counting of degrees of freedom (d.o.f.). The double-traceless field $h$ has
$N_{h}=\binom{D+s-1}{D-1}-\binom{D+s-5}{D-1}$ (4.17)
@@
$N_{\Lambda}=\binom{D+s-2}{D-1}-\binom{D+s-4}{D-1}-\left[\binom{D+s-4}{D-1}-\binom{D+s-6}{D-1}\right$]. (4.18)
The equations of motion lead to the gauge independent constraints (4.13) and (4.14) which are traceless rank-(s-3) symmetric tensors. The number of independent constraints is
$N_{\text{constraints}}=2\left[\binom{D+s-4}{D-1}-\binom{D+s-6}{D-1}\right$] (4.19)
@@
$N_{dof}$ $=$ $N_{h}-2N_{\Lambda}-N_{\text{constraints}}$ (4.20)
@@
For arbitrary $D$ dimensions and arbitrary spin-$s$, (4.20) coincides exactly with the number of d.o.f. of the Fronsdal model. The inclusion of higher derivative Lagrangian constraints in the counting of degrees of freedom is typical of gauge parameters with derivative restrictions, see chapter 5 of [8]. This is the case also of the spin-2 TDiff model (3.2) whose equations of motion lead to the two derivative scalar constraint $\partial^{\alpha}\partial^{\beta}h_{\alpha\beta}=0$. So, although we have (in $D=4$) only 3 independent gauge parameters due to the transverse condition $\partial^{\mu}\epsilon^{T}_{\mu}=0$, we can fix the four gauge conditions $f_{\mu}=\partial^{\mu}h_{\mu\nu}=0$ since they are on shell transverse due to the two derivative scalar constraint. Taking into account the “regauging” we have $N_{dof}=10-2\times 3-1=3$ in agreement with one spin-0 and the $\pm 2$ helicities of the spin-2 particle [^8] Remarkably, the Hamiltonian analysis of [19] has shown that the spin-2 TDiff model has 2D12D-1, instead of the expected 2(D1)2(D-1), first class constraints and no second class constraints. The Hamiltonian constraints analysis of the PBF model for arbitrary integer spin-$s$ is in progress..
@@ 4.3.1 The spin-3 case
$\delta h_{\mu\nu\alpha}=ip_{(\mu}\overline{\Lambda}^{\ddagger}_{\nu\alpha)}.$ (4.21)
@@
$p\!\cdot\!p\!\cdot\!\overline{\Lambda}^{\ddagger}=0=\eta\cdot\overline{\Lambda}^{\ddagger}.$ (4.22)
Assuming $p_{+}\neq 0$, due to (4.22), the gauge parameters $\overline{\Lambda}^{\ddagger}_{--}$ and $\overline{\Lambda}^{\ddagger}_{+-}$ become functions of $(\overline{\Lambda}^{\ddagger}_{++},\overline{\Lambda}^{\ddagger}_{IJ})$. However, we can use the whole set of unconstrained parameters $(\overline{\Lambda}^{\ddagger}_{++},\overline{\Lambda}^{\ddagger}_{IJ},\overline{\Lambda}^{\ddagger}_{+J},\overline{\Lambda}^{\ddagger}_{-J})$ in order to impose the maximal number of gauge conditions:
$h_{+++}=0=h_{+IJ}=h_{++J}=h_{+-J}\,.$ (4.23)
@@
$E_{\mu\nu\rho}(p)\equiv p^{2}h_{\mu\nu\alpha}-p_{(\mu}p^{\rho}h_{\nu\alpha)\rho}+p_{(\mu}p_{\nu}h_{\alpha)}=0.$ (4.24)
Henceforth we use the short hand notation: $(\mu,\nu,\rho)\equiv E_{\mu\nu\rho}(p)$ and assume that (4.23) is holding true. From $(+,+,+)=0$ one finds $h_{-++}=0$. The equations $(+,+,J)=0$ and $(+,I,J)=0$ lead to $h_{IIJ}=0$ and $h_{-IJ}=\frac{1}{p^{+}}p_{K}h_{IJK}$, respectively, they lead altogether to $h_{-JJ}=0$. The constraint $\partial\!\cdot\!h^{\prime}=0$ implies $h_{+--}=0$. The field is thus on shell traceless ($h_{\alpha}=-2\,h_{+-\alpha}+h_{II\alpha}=0$) and satisfies $h_{+\mu\nu}=0$. From $(+,\nu,\alpha)=0$ we have the transverse constraint $p^{\rho}h_{\rho\nu\alpha}=0$ which at $(\nu,\alpha)=(-,k)$ and $(\nu,\alpha)=(-,-)$ leads, using previous results, to
$h_{--K}=\frac{p_{I}\,p_{J}\,h_{IJK}}{p_{+}^{2}}\quad;\quad h_{---}=\frac{p_{I}\,p_{J}\,p_{K}h_{IJK}}{p_{+}^{3}}.$ (4.25)
So we can consider $h_{IJK}$ as the only relevant field components. With $h_{\alpha}=0=p^{\rho}h_{\rho\mu\nu}$ back in (4.24) we have $p^{2}h_{\mu\nu\alpha}=0$. Effectively, we have the following physical equations:
$p^{2}h_{IJK}=0,\quad h_{JJK}=0.$ (4.26)
@@
$p^{2}h_{\mu\nu\alpha\beta}-p_{(\mu}\ p^{\rho}h_{\rho\nu\alpha\beta)}+p_{(\mu}\ p_{\nu}\ h^{\prime}_{\alpha\beta)}\equiv(\mu,\nu,\alpha,\beta)=0.$ (4.28)
@@
$h_{KKIJ}=0.$ (4.29)
From (+++)=0(-+++)=0 and the constraint $(p\cdot h^{{}^{\prime}})_{+}=0$, we have $h_{++--}=0=h_{+-KK}$. Now we make use of the only unconstrained gauge parameters left over. Namely, since $\delta h_{+-IJ}=i\,p_{(+}\overline{\Lambda}^{\ddagger}_{IJ-)}$ and we already have $h_{+-KK}=0$, we only need the space traceless parameter $\overline{\Lambda}^{\ddagger}_{\overline{I}\overline{J}-}=\overline{\Lambda}^{\ddagger}_{IJ-}-\delta_{IJ}\overline{\Lambda}^{\ddagger}_{KK-}/(D-2)$ in order to fix our last light cone gauge condition: $h_{+-\overline{I}\overline{J}}=0$. Consequently, we have $h_{+-IJ}=0$ henceforth. There is no remaining unconstrained parameter anymore, since $\overline{\Lambda}^{\ddagger}_{KK-}$ is already constrained as we have argued before (4.27).
@@
$h_{---I}=\frac{p_{J}\,p_{K}\,p_{L}\,h_{IJKL}}{p_{+}^{3}}\quad$ ; $h_{----}=\frac{p_{I}\,p_{J}\,p_{K}\,p_{L}\,h_{IJKL}}{p_{+}^{3}}$ (4.30)
@@
$p^{2}h_{IJKL}=0.$ (4.31)
Altogether with (4.29) we have (D+4)(D3)(D2)(D1)/(4!)(D+4)(D-3)(D-2)(D-1)/(4\ !) massless propagating modes as in the Fronsdal spin-4 model.
@@ 4.3.3 Arbitrary integer spin
$h_{+(k)J(s-k)}=0\quad,\quad k=1,\cdots,s\,.$ (4.32)
@@
$p^{2}\,h_{\mu_{1}\cdots\mu_{s}}+p_{(\mu_{1}}p_{\mu_{2}}\,h^{{}^{\prime}}_{\mu_{3}\,\mu_{4}\cdots\mu_{s})}-p_{(\mu_{1}}p^{\rho}h_{\rho\,\mu_{2}\cdots\mu_{s})}\equiv\left(\mu_{1}\,\mu_{2}\cdots\mu_{s}\right)=0.$ (4.33)
@@
$h_{++\mu(s-2)}=0\quad\,\text{and}\,\quad h_{+KK\mu(s-3)}=0,$ (4.35)
@@
$h_{-(l)I(s-l)}=\frac{1}{(p_{+})^{l}}p_{I_{1}}p_{I_{2}}\dots p_{I_{l}}h_{I_{1}I_{2}\dots I_{l}I(s-l)},$ (4.37)
where $l\in\{1,2,\dots,s\}$. Notice that the constraint $p\cdot p\cdot p\cdot h=0$ is equivalent to (4.37). So the only nontrivial field components are given by $h_{J(s)}$. The equations of motion $\left(I_{1}\cdots I_{s}\right)=0$ become simply
$p^{2}h_{I(s)}=0.$ (4.38)
Since we still have the algebraic constraints (4.34) we end up with two helicities in $D=4$. The condition on the gauge parameter $\partial\cdot\partial\cdot\overline{\Lambda}^{\ddagger}=0$ is compensated by the constraint $\partial\cdot h^{{}^{\prime}}=0$. Notice that they are both totally symmetric rank-(s-3) traceless tensors.
@@ 5 Conclusion and perspectives
We have found here a partially broken Fronsdal (PBF) model for arbitrary integer spin, it is defined in (4.8), (4.7) and (4.1). The model is invariant under traceless higher spin diffeomorphisms with vanishing double divergence, see (4.11). In subsection 3.3 we have verified the content of the spin-3 PBF model in a gauge invariant way while in subsection 4.3 we have checked the spectrum of the arbitrary integer spin case with $s\geq 3$ in the light cone gauge. We conclude that the PBF model has the same propagating content of the Fronsdal model, i.e., only the $\pm s$ helicities in $D=4$.
@@
$\partial\cdot T=\eta\,\Omega+\partial\,\partial\,J,$ (5.1)
where $\Omega$ and $J$ are rank-(s-3) symmetric tensors constrained only by the fact that the double trace of the right hand side of (5.1) must vanish for $s\geq 5$. In the Fronsdal theory the non conservation law is more restrictive ($J=0$).
@@
Regarding a curved background, since $\nabla\cdot h^{{}^{\prime}}$ is by itself invariant under $\delta h=\nabla\overline{\Lambda}^{\ddagger}$, with $g^{\mu_{1}\mu_{2}}\,\overline{\Lambda}^{\ddagger}_{\mu_{1}\mu_{2}\cdots\mu_{s-1}}=0=\nabla\cdot\nabla\cdot\overline{\Lambda}^{\ddagger}$, it follows that the same non minimal compensating terms of the Fronsdal action [25] on (A)dS backgrounds, quadratic on $h$ and $h^{{}^{\prime}}$, will turn the model $\pazocal{L}^{(s)}(1,1,c)$ given in (4.2), with $\partial_{\mu}\to\nabla_{\mu}$, invariant under gauge transformations on (A)dS.
@@
It is important to mention that the PBF model is a two parameters theory, $\pazocal{L}_{PBF}^{(s)}=\pazocal{L}^{(s)}[a,b_{S}^{(s)}(a,c),c]=\pazocal{L}_{PBF}^{(s)}(a,c)$. We can always redefine one of the parameters as we wish by means of local field redefinitions that we have named r-shifts, see (4.2) and (4.3). We are currently investigating the role of the remaining parameter which can not be shifted, along the lines of [21] where we have shown, inspired also by non-local field redefinitions, that the two parameters TDiff spin-3 model $S[a,b,c_{S}(a,b)]$, see (3.5), is physically equivalent to its simplest version $S(0,0,0)$ which corresponds to the spin-3 case of the Maxwell-like model of [22]. Similarly, we believe that non-local field redefinitions may help to clarify the role of the left over constant parameter of the PBF model.